Hello
I have a question on my homework about inverse relations and functions. this is the question.
Find the inverse of y=sqrtx+9 Is the inverse a function?
I think i found the 1st part which is y=x^2-9 but i dont know if the inverse is a function?
Thanks for the help!

Answer by stanbon(48535)

Find the inverse of y=sqrt(x+9)
Note: this function is defined for x>=-9
----------
Inverse:
x = sqrt(y+9)
x^2 = y+9
y = x^2-9
--------------
Note: y= x^2-9 is a function. As an inverse to
y=sqrt(x+9) it is a function defined for x>=0
====================
Cheers,
Stan H.